OpenStudy (helpblahblahblah):

Help please What is the approximate equation of this line of best fit in slope-intercept form? A. y = −13x + 13 over 10 B. y = negative 13 over 10x + 13 C. y = negative 10 over 13x + 13 D. y = −13x + 10 over 13

OpenStudy (helpblahblahblah):

@TwilightRain1

OpenStudy (anonymous):

was da question?

OpenStudy (helpblahblahblah):

What is the approximate equation of this line of best fit in slope-intercept form?

OpenStudy (anonymous):

omg, i dunno :P

OpenStudy (helpblahblahblah):

Lol okay @parakeet @EclipsedStar

OpenStudy (helpblahblahblah):

@sleepyjess

OpenStudy (perl):

you can pick two nice points to get the equation of that line,

OpenStudy (helpblahblahblah):

I've been trying to do that but I kinda forgot how to .-.

OpenStudy (helpblahblahblah):

Okay

OpenStudy (perl):

lets try the points (0,13) and (7,4)

OpenStudy (perl):

we can use slope formula m = (4-13) / (7 - 0) = -9/7

OpenStudy (helpblahblahblah):

Okay

OpenStudy (perl):

and the y intercept is 13

OpenStudy (helpblahblahblah):

So its in between B and C

OpenStudy (perl):

so one equation for this line looks like y = -9/7x + 13 but thats not one of the given choices. but notice that -9/7 is close to -13/10 , decimal value

OpenStudy (perl):

-9/7 = -1.29 -13/10 = -1.3

OpenStudy (helpblahblahblah):

So B?